OpenStudy (anonymous):

what is the slope of the line with equation 6x-10y=8?

OpenStudy (anonymous):

The Slope would be 6/10 because ... 6x-10y = 8 add 10y to both sides 6x = 10y + 8 subtract 8 from both sides 6x-8 = 10y divide each side by 10 (6/10)x - (8/10) = y 6/10 is the slope

OpenStudy (campbell_st):

simplify the fraction for the slope of the line

OpenStudy (anonymous):

simplified would be 3/5

OpenStudy (anonymous):

make y the subject of the formular! y=10/6x-8/10 its in the form: y=mx+c, where m is the slope/gradient of the line and c is the point where the line crosses the y-axis